# Quantile Reward Policy Optimization: Alignment with Pointwise Regression and Exact Partition Functions

### v1

@@ Abstract
Method Avg. Pass (%$\uparrow$
Llama 8B Tülu 3 SFT ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}20.9}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.5}}$
+ SFT on Correct ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}18.8}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}1.1}}$
+ REBEL ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}26.1}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}1.8}}$
+ DPO ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}30.2}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}1.4}}$
+ SimPO ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}22.3}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}1.4}}$
+ QRPO (Ours) ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}\mathbf{32.7}}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}1.0}}$
@@ 1 Introduction
Alignment methods are highly effective in fine-tuning large language models (LLMs), improving their conversational abilities (Ouyang et al., 2022; Stiennon et al., 2020; Rafailov et al., 2023), safety (Bai et al., 2022), and reasoning capabilities (Lambert et al., 2024; Shao et al., 2024a; Guo et al., 2025). They are applied as the final stage of the LLM training pipeline after pre-training the LLM on a large corpus of data (Brown et al., 2020) and fine-tuning the LLM to follow instructions (Ouyang et al., 2022). However, while simple and efficient methods, such as DPO (Rafailov et al., 2023), relying on a preference reward signal, achieve strong performance in conversational tasks, more complex and computationally intensive methods, such as GRPO (Shao et al., 2024b), have been the only alternative to effectively optimize a pointwise reward signal.
@@
Alignment methods are often referred to by different names, such as reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), reinforcement learning from verifiable rewards (RLVR) (Lambert et al., 2024), and direct alignment algorithms (DAA) (Rafailov et al., 2024), but the underlying algorithms in most of these methods optimize a common objective from regularized reinforcement learning (RL) (Jaques et al., 2017). This nomenclature essentially distinguishes two dimensions: whether the reward is optimized explicitly (RLHF, RLVR) or implicitly (DAA) and whether it is modeled from (human) preference feedback (RLHF, DAA) or objectively defined (RLVR). However, another crucial dimension is the data regime required by the underlying algorithm, which significantly impacts the computational cost of the methods. Policy improvement (PI) methods, such as PPO (Ouyang et al., 2022; Schulman et al., 2017) and GRPO (Shao et al., 2024b), require online sampling as they perform online policy improvement, thus increasing the complexity and cost of distributing large networks across training and inference engines. In contrast, policy fitting (PF) methods, such as DPO (Rafailov et al., 2023) and REBEL (Gao et al., 2024), can leverage any data distribution because they aim to fit a closed-form solution of the regularized RL objective valid for all data, giving them the reputation for being simpler and more stable, as usually used with offline data.111 In the LLM literature, the terms are “RL” methods and “DAA” methods, but we feel that this is not accurate as all methods optimize an RL objective, and methods like REBEL do not fit within the “direct” naming which refers to directly using a preference model, which is not necessary to target the closed-form solution optimization.
@@
Existing PF methods rely on relative rewards, i.e., differences between rewards, to tractably fit the closed-form solution, as estimating it for absolute rewards is known to be intractable due to the difficulty of estimating the partition function. This feature has contributed to their popularity, as human feedback can be expressed as preferences and modeled as relative rewards by preference models. However, relative rewards can provide a suboptimal signal, for example, when both the chosen and rejected completions are better than the training model. This leads to conflicting dynamics with updates from samples that could both improve the model. Moreover, human feedback can be more easily collected as absolute feedback in the form of a rating scale. More pressingly, with the emergence of powerful absolute reward models for human alignment trained with regression such as ArmoRM (Wang et al., 2024) and Nemotron-Reward (Wang et al., 2025) and successful absolute-reward applications such as verifiable rewards (Lambert et al., 2024), PF methods are losing applicability and practitioners have to resort to complex PI methods to use those rewards.
@@
    QRPO consistently achieves top performance in conversational and coding tasks We show through an extensive and fair experimental protocol that QRPO consistently obtains the best rewards and AlpacaEval (Li et al., 2023) scores. We compared our method to DPO, REBEL, and SimPO, where each was trained in more than 200 settings of models, datasets, hyperparameters, and distribution shifts, including Llama 8B and Mistral 7B as models and Magpie-Air (Xu et al., 2024b) and UltraFeedback (Cui et al., 2024) as datasets. We also demonstrate that QRPO is effective in a LeetCode (Xia et al., 2025) coding task, where the reward is canonically pointwise, expressing the test-case pass rate and loses signal when expressed as a preference.
@@ 2 Preliminaries
The underlying RL fine-tuning objective in alignment methods is formulated as (Jaques et al., 2017; Stiennon et al., 2020; Rafailov et al., 2023)
@@
PPO (Schulman et al., 2017) and their variants rewrite the RL fine-tuning objective by moving the KL penalty to the reward (Stiennon et al., 2020, PPO), or consider it as an auxiliary loss (Shao et al., 2024b, GRPO), to then effectively optimize a standard RL objective by taking policy gradient steps to improve the policy online. Hence, these methods require sampling before each gradient step, requiring extra complexity to host inference instances of the policy, and making them more costly in terms of time and computational resources.
@@
Policy fitting methods leverage this expression to fit the optimal policy. By substituting the optimal policy $\pi^{*}$ with a parameterized model $\pi_{\theta}$ in Equation 2, we can formulate a distribution matching or regression task. However, estimating $\pi^{*}$ directly is known to be intractable due to the difficulty in estimating the partition function $Z$ (Rafailov et al., 2023). Hence, current methods resort to relative signals between two completions to eliminate the absolute partition function. Indeed, we have:
$\mathcal{R}(x,y)=\beta\log{\frac{\pi^{*}(y\mid x)}{\pi_{ref}(y\mid x)}}+\beta\log{Z(x)},\vspace{-.2\baselineskip}$ (3)
@@
When the reward is used to parameterize a preference model, such as the Bradley-Terry model (Bradley & Terry, 1952), this difference gives the preference between two completions $p(y^{+}\succ y^{-})=\sigma(\mathcal{R}(x,y^{+})-\mathcal{R}(x,y^{-}))$, and direct preference optimization methods such as DPO (Rafailov et al., 2023) learn the policy by substituting the reward difference by the policy difference and fitting the preferences. Gao et al. (2024) propose REBEL, which fits the optimal policy by directly regressing the difference in $\log$-probability ratios to the difference in rewards (as shown in Equation 4). The REBEL algorithm was originally presented as an iterative method; however, in this work, we focus on the RL fine-tuning objective with a fixed reference policy, and as the first iteration of REBEL already fits the optimal policy relative to $\pi_{ref}$ we present REBEL as a policy fitting method.
@@ 3 Quantile Regression Policy Optimization (QRPO)
We choose the mean-squared error (MSE) as a loss, which is also a common choice in other LLM alignment algorithms (Azar et al., 2024). Moreover, as a consequence of the central limit theorem, the noise at the target value is likely to have a Gaussian distribution, which makes the MSE appropriate.
The objective in Equation 6 is the most straightforward regression objective to derive from Equation 2 and is aimed at by previous work (Gao et al., 2024; Pierre Harvey Richemond et al., 2024), but is still intractable due to the difficulty in estimating the partition function $Z$, leading to these previous attempts to resort to approximations (Pierre Harvey Richemond et al., 2024) or to forgo it entirely (Gao et al., 2024). However, this form has multiple benefits, and QRPO’s main contribution is to transform the reward to maintain this regression objective while making the partition function tractable.
@@
First, unlike policy improvement methods such as PPO, the training completion $y$ does not have to be online with respect to $\pi_{\theta}$ and can be sampled from any distribution, such as an offline dataset to leverage pre-existing data (as in DPO originally), online data when novel exploration is needed, as in Online DPO (Qi et al., 2024), or a mix of offline, off-policy, and on-policy data. This is a significant benefit of policy fitting methods.
Second, the primary rationale for employing an absolute reward signal instead of a relative one (as in DPO or REBEL) is that an absolute signal not only contains strictly more information but also qualitatively influences the optimization dynamics when data coverage is limited. Relative reward methods leverage limited data coverage by increasing the difference between paired samples, often simultaneously lowering the probabilities of both (Pal et al., 2024). This consistent reduction in probabilities leads to unintended probability mass drifting toward out-of-distribution samples, resulting in unstable or ineffective long-term training (Razin et al., 2025). In contrast, an absolute reward approach enforces model probabilities to match fixed, predefined targets, directing probability mass explicitly toward higher-reward samples rather than drifting towards out-of-distribution outcomes, enabling stable and effective optimization. See Appendix A for a comprehensive discussion.
@@
The main argument commonly supporting the difficulty of estimating $Z(x)$ in Equations 6 and 2 is the intractability of the infinite sum over completions $y$ for an LLM. Our first insight is to show that the partition function can be expressed using only the distribution of reference rewards $\mathcal{R}(x,y)$ for completions $y$ drawn from the reference policy $\pi_{ref}(\cdot\mid x)$. Let $P_{ref}(r|x)$ be the probability of a reference reward $r$, and $M$ denote the moment generating function (MGF). We show that (proof in Appendix C):
@@
This derivation allows us to have a better intuition of the challenge of estimating the partition function: it is particularly challenging with reward distributions that have a right tail because the exponent of large rewards, which happen with low probability, makes the estimate very noisy. We support this intuition with a theoretical analysis of the noise amplitude of $Z(x)$ when rewards follow a standard normal distribution, which can be an example for typical LLM reward distributions, such as in our experiments, which do have a right tail. We show that with $\beta=0.1$, approximately $10^{40}$ samples are required to achieve a reasonable signal-to-noise ratio in the estimate of $Z(x)$ (proof in Appendix G).
@@
When the reference reward distribution can be treated as continuous, such as with a reward model or a coding test-case pass rate, as in our experiments, the resulting distribution of the quantile-transformed reference rewards $\mathbf{\mathcal{R}}_{q}(x,y)$ for $y\sim\pi_{ref}(\cdot\mid x)$ is uniform, independently of the reference policy (Appendix D).222 For completeness, we also derive QRPO for distributions with a small number of possible values in App. E by directly estimating the expectation for quantile rewards, as the quantile does not result in a known distribution. Therefore, using Eq. 7 which holds for any rewards, the corresponding partition function for this new reward is simply the MGF of the uniform distribution (details in Appendix E):
@@
For computational stability, we neglect the $-1$ term as $\exp(1/{\beta})$ is typically very large for useful values of $\beta\leq 0.1$, and use $\log(\beta)+1/{\beta}$ for $\log Z$, obtaining
@@
We can introduce a general transformation $f$ on top of the quantile reward that still allows us to analytically derive the corresponding partition function in two different ways when the reference rewards can be treated as continuous: using the MGF of this new reward if the transformation leads to a known distribution or computing the following integral (proof in App. E):
[…]
where $f$ represents any function defined on the interval $[0,1]$. This formulation demonstrates the significant flexibility of QRPO, positioning it as a general framework and allowing the design of reward distributions with specific properties that shape the structure of the optimal policy.
Invariance of the optimal policy
We perform an initial analysis of the effect of transformation functions in QRPO: we show in Appendix J that for a class of functions $f(t)$ that are strictly increasing and finite at $t=1$ the resulting optimal policy and thus the QRPO result is invariant to the specific choice of $f$ given a small enough $\beta$. However, preliminary experiments on our tasks suggest that selecting $f(t)=\log{t}$, as in the vBoN and BOND algorithms (Amini et al., 2025; Sessa et al., 2025), slightly reduces QRPO’s performance. Indeed, in practice, as the optimal policy is hard to achieve, different transformations can result in different final performances due to the optimization dynamics they induce. Hence, studying transformation functions in QRPO is a key direction for future work.
@@
The optimal policy of the RL fine-tuning objective depends on the reward and is thus different when we transform the reward. Using results from Amini et al. (2025) and Sessa et al. (2025) showing that optimizing the $\log$-quantile reward in the RL fine-tuning objective induces a Best-of-$N$ policy, and our invariance result discussed above, we can interpret the QRPO optimal policy as approximating a Best-of-$N$ policy for $N=1/\beta+1$ (Sessa et al., 2025). Its reward distribution when the reference reward distribution is Gaussian is shown in Figure 5 and Appendix B describes how to choose the transformation $f$ to shape the optimal policy.
@@
We have eliminated the noise in the partition function by using an exact expression computed analytically. However, we have introduced new noise in the regression objective, originating from the transformed reward $\mathbf{\mathcal{R}}_{q}(x,y)$, as it requires estimating a quantile. We show in Appendix G, using a similar analysis as in Insight 1 with a Gaussian distribution, that the noise in the calibrated target (Eq. 6) computed by keeping the original reward and estimating $Z(x)$ is significantly higher than when estimating the quantile reward and using the exact expression of $Z(x)$, in particular for the order of magnitude of $\beta$ commonly used in RL fine-tuning ($0.1$ and lower). Hence, QRPO reduces the noise in the objective function dramatically, and we therefore enable the general approach of policy fitting with pointwise absolute rewards to work well in practice.
@@
$\mathbb{E}_{(x_{i},y_{i})\sim\mathcal{D}}\left[\mathcal{R}_{q}(x_{i},y_{i})-\beta\log{\beta}-1-\beta\log{\frac{\pi_{\theta}(y_{i}\mid x_{i})}{\pi_{ref}(y_{i}\mid x_{i})}}\right]^{2}\vspace{-.4\baselineskip}$

### v2

@@ Abstract
[…]
@@ 1 Introduction
Alignment methods are highly effective in fine-tuning large language models (LLMs), improving their conversational abilities (ouyang_training_2022; stiennon2020learning; rafailov2023direct), safety (bai2022training), and reasoning capabilities (lambert2024t; shao2024deepseekmath; guo2025deepseek). They are applied as the final stage of the LLM training pipeline after pre-training the LLM on a large corpus of data (brown2020language) and fine-tuning the LLM to follow instructions (ouyang_training_2022). However, while simple and efficient methods, such as DPO (rafailov2023direct), relying on a preference reward signal, achieve strong performance in conversational tasks, more complex and computationally intensive methods, such as GRPO (shao2024grpo), have been the only alternative to effectively optimize a pointwise reward signal.
@@
Alignment methods are often referred to by different names, such as reinforcement learning from human feedback (RLHF) (ouyang_training_2022), reinforcement learning from verifiable rewards (RLVR) (lambert2024t), and direct alignment algorithms (DAA) (rafailov2024scaling), but the underlying algorithms in most of these methods optimize a common objective from regularized reinforcement learning (RL) (jaques2017sequence). This nomenclature distinguishes two dimensions: whether the reward is optimized explicitly (RLHF, RLVR) or implicitly (DAA) and whether it is modeled from (human) preference feedback (RLHF, DAA) or objectively defined (RLVR). Yet, another crucial dimension is the data regime required by the underlying algorithm, which significantly impacts the computational cost of the methods. Policy improvement (PI) methods, such as PPO (ouyang_training_2022; schulman_proximal_2017), GRPO (shao2024grpo), and RLOO (ahmadian2024rloo), require online sampling as they perform online policy improvement, thus increasing the complexity and cost of distributing large networks across training and inference engines. In contrast, policy fitting (PF) methods, such as DPO (rafailov2023direct) and REBEL (gao_rebel_2024), can leverage any data distribution because they aim to fit a closed-form solution of the regularized RL objective valid for all data, giving them the reputation for being simpler and more stable, as usually used with offline data. [^1] In the LLM literature, the terms are “RL” methods and “DAA” methods, but we feel that this is not accurate as all methods optimize an RL objective, and methods like REBEL do not fit within the “direct” naming which refers to directly using a preference model, which is not necessary to target the closed-form solution optimization.
@@
Existing PF methods rely on relative rewards, i.e., differences between rewards, to tractably fit the closed-form solution, as estimating it for absolute rewards is known to be intractable due to the difficulty of estimating the partition function. This feature has contributed to their popularity, as human feedback can be expressed as preferences and modeled as relative rewards by preference models. However, relative rewards can provide a suboptimal signal, for example, when both the chosen and rejected completions are better than the training model. This leads to conflicting dynamics with updates from samples that could both improve the model. Moreover, human feedback can be more easily collected as absolute feedback in the form of a rating scale. More pressingly, with the emergence of powerful absolute reward models for human alignment trained with regression such as ArmoRM (wang_interpretable_2024) and Nemotron-Reward (wanghelpsteer2) and successful absolute-reward applications such as verifiable rewards (lambert2024t), PF methods are losing applicability and practitioners have to resort to complex PI methods to use those rewards.
@@
    QRPO consistently achieves top performance in conversational and coding tasks We show through an extensive and fair experimental protocol that QRPO consistently obtains the best rewards and AlpacaEval (alpaca_eval) scores. We compared our method to DPO, REBEL, and SimPO, where each was trained in more than 200 settings of models, datasets, hyperparameters, and distribution shifts, including Llama 8B and Mistral 7B as models and Magpie-Air (xu_magpie_2024) and UltraFeedback (cui_ultrafeedback_2024) as datasets. We also demonstrate that QRPO is effective in a LeetCode (xia2025leetcodedataset) coding task, where the reward is canonically pointwise, expressing the test-case pass rate and loses signal when expressed as a preference.
@@ 2 Preliminaries
The underlying RL fine-tuning objective in alignment methods is formulated as (jaques2017sequence; stiennon2020learning; rafailov2023direct)
@@
PPO (schulman_proximal_2017) and their variants rewrite the RL fine-tuning objective by moving the KL penalty to the reward (stiennon2020learning, PPO), or consider it as an auxiliary loss (shao2024grpo, GRPO), to then effectively optimize a standard RL objective by taking policy gradient steps to improve the policy online. Hence, these methods require sampling before each gradient step, requiring extra complexity to host inference instances of the policy, and making them more costly in terms of time and computational resources.
@@
Policy fitting methods leverage this expression to fit the optimal policy. By substituting the optimal policy $\pi^{*}$ with a parameterized model $\pi_{\theta}$ in Equation 2, we can formulate a distribution matching or regression task. However, estimating $\pi^{*}$ directly is known to be intractable due to the difficulty in estimating the partition function $Z$ (rafailov2023direct). Hence, current methods resort to relative signals between two completions to eliminate the absolute partition function. Indeed, we have:
$\mathcal{R}(x,y)=\beta\log{\frac{\pi^{*}(y\mid x)}{\pi_{ref}(y\mid x)}}+\beta\log{Z(x)},\vskip-2.39996pt$ (3)
@@
When the reward is used to parameterize a preference model, such as the Bradley-Terry model (bradley1952rank), this difference gives the preference between two completions $p(y^{+}\succ y^{-})=\sigma(\mathcal{R}(x,y^{+})-\mathcal{R}(x,y^{-}))$, and direct preference optimization methods such as DPO (rafailov2023direct) learn the policy by substituting the reward difference by the policy difference and fitting the preferences. gao_rebel_2024 propose REBEL, which fits the optimal policy by directly regressing the difference in $\log$-probability ratios to the difference in rewards (as shown in Equation 4). The REBEL algorithm was originally presented as an iterative method; however, in this work, we focus on the RL fine-tuning objective with a fixed reference policy, and as the first iteration of REBEL already fits the optimal policy relative to $\pi_{ref}$ we present REBEL as a policy fitting method.
@@ 3 Quantile Regression Policy Optimization (QRPO)
We choose the mean-squared error (MSE) as a loss, which is also a common choice in other LLM alignment algorithms (azar2023ipo). Moreover, as a consequence of the central limit theorem, the noise at the target value is likely to have a Gaussian distribution, which makes the MSE appropriate.
The objective in Equation 6 is the most straightforward regression objective to derive from Equation 2 and is aimed at by previous work (gao_rebel_2024; pierre_harvey_richemond_offline_2024), but is still intractable due to the difficulty in estimating the partition function $Z$, leading to these previous attempts to resort to approximations (pierre_harvey_richemond_offline_2024) or to forgo it entirely (gao_rebel_2024). However, this form has multiple benefits, and QRPO’s main contribution is to transform the reward to maintain this regression objective while making the partition function tractable.
@@
First, unlike policy improvement methods such as PPO, the training completion $y$ does not have to be online with respect to $\pi_{\theta}$ and can be sampled from any distribution, such as an offline dataset to leverage pre-existing data (as in DPO originally), online data when novel exploration is needed, as in Online DPO (qi2024online), or a mix of offline, off-policy, and on-policy data. This is a significant benefit of policy fitting methods.
Second, the primary rationale for employing an absolute reward signal instead of a relative one (as in DPO or REBEL) is that an absolute signal not only contains strictly more information but also qualitatively influences the optimization dynamics when data coverage is limited. Relative reward methods leverage limited data coverage by increasing the difference between paired samples, often simultaneously lowering the probabilities of both (pal2024smaugfixingfailuremodes). This consistent reduction in probabilities leads to unintended probability mass drifting toward out-of-distribution samples, resulting in unstable or ineffective long-term training (razin2025unintentional). In contrast, an absolute reward approach enforces model probabilities to match fixed, predefined targets, directing probability mass explicitly toward higher-reward samples rather than drifting towards out-of-distribution outcomes, enabling stable and effective optimization. See Appendix A for a comprehensive discussion.
@@
The main argument commonly supporting the difficulty of estimating $Z(x)$ in Equations 6 and 2 is the intractability of the infinite sum over completions $y$ for an LLM. Our first insight is to show that the partition function can be expressed using only the distribution of reference rewards $\mathcal{R}(x,y)$ for completions $y$ drawn from the reference policy $\pi_{ref}(\cdot\mid x)$. Let $P_{ref}(r|x)$ be the probability of a reference reward $r$, and $M$ denote the moment generating function (MGF). We show that (proof in Appendix D):
@@
This derivation allows us to have a better intuition of the challenge of estimating the partition function: it is particularly challenging with reward distributions that have a right tail because the exponent of large rewards, which happen with low probability, makes the estimate very noisy. We support this intuition with a theoretical analysis of the noise amplitude of $Z(x)$ when rewards follow a standard normal distribution, which can be an example for typical LLM reward distributions, such as in our experiments, which do have a right tail. We show that with $\beta=0.1$, approximately $10^{40}$ samples are required to achieve a reasonable signal-to-noise ratio in the estimate of $Z(x)$ (proof in Appendix H).
@@
When the reference reward distribution can be treated as continuous, such as with a reward model or a coding test-case pass rate, as in our experiments, the resulting distribution of the quantile-transformed reference rewards $\mathbf{\mathcal{R}}_{q}(x,y)$ for $y\sim\pi_{ref}(\cdot\mid x)$ is uniform, independently of the reference policy (Appendix E). [^2] For completeness, we also derive QRPO for reward distributions with a small number of possible values in App. F by directly estimating the expectation for quantile rewards, as they do not result in a known distribution. Therefore, using Eq. 7 which holds for any rewards, the corresponding partition function for this new reward is simply the MGF of the uniform distribution (details in Appendix F):
@@
For computational stability, we neglect the 1-1 term as $\exp(1/{\beta})$ is typically very large for useful values of $\beta\leq 0.1$, and use $\log(\beta)+1/{\beta}$ for $\log Z$, obtaining
@@
We can generalize the derivation of QRPO to more general reward transformations. Specifically, we can apply a wide range of functions $f$ on top of the quantile reward, while still retaining the ability to analytically derive the corresponding partition function when the reference rewards can be treated as continuous (proof in App. F):
$\mathbf{\tilde{\mathcal{R}}}(x,y)=f\left(\mathbf{\mathcal{R}}_{q}(x,y)\right);\ \ \ \ \tilde{Z}(x)=M_{\tilde{r}}\left(1/{\beta}\right)=\int_{0}^{1}\exp\Bigl(\tfrac{1}{\beta}f(t)\Bigr)\,dt.\vskip-3.60004pt$ (13)
where $f$ represents any function defined on the interval $[0,1]$. This formulation demonstrates a significant flexibility of QRPO, positioning it as a general framework and allowing the design of reward distributions with specific properties that shape the structure of the optimal policy.
@@
The optimal policy of the RL fine-tuning objective depends on the reward. Therefore, optimizing the quantile reward (Eq. 9) or its generalized transformed version (Eq. 13) produces policies with different properties compared to optimizing the original reward. Notably, balashankar2024infalign demonstrate that optimizing the quantile reward is equivalent to optimizing the win rate against the reference policy and further allows to optimize the policy performance for a specific inference-time scaling strategy by choosing a specific transformation $f$ on top of the quantile reward in Eq. 13. This highlights a key advantage of the QRPO framework.
Foundations for theoretical analyses
The analytical tractability of the partition function in QRPO is just one example of the broader opportunities for theoretical analysis enabled by quantile-based reward framework. To illustrate this, we conduct an initial investigation into the impact of different transformation functions in Eq. 13. In Appendix K, we show that for a class of functions $f(t)$ that are strictly increasing and finite at $t=1$, the resulting optimal policy is invariant to the specific choice of $f$ when $\beta$ is sufficiently small. However, understanding the effects of other types of functions, the relationship between the optimal policy and the optimization dynamics, and several other related questions remain open problems and represent key directions for future research.
@@
We have eliminated the noise in the partition function by using an exact expression computed analytically. However, we have introduced new noise in the regression objective, originating from the transformed reward $\mathbf{\mathcal{R}}_{q}(x,y)$, as it requires estimating a quantile. We show in Appendix H, using a similar analysis as in Insight 1 with a Gaussian distribution, that the noise in the calibrated target (Eq. 6) computed by keeping the original reward and estimating $Z(x)$ is significantly higher than when estimating the quantile reward and using the exact expression of $Z(x)$, in particular for the order of magnitude of $\beta$ commonly used in RL fine-tuning ($0.1$ and lower). Hence, QRPO reduces the noise in the objective function dramatically, and we therefore enable the general approach of policy fitting with pointwise absolute rewards to work well in practice.
@@
$\mathbb{E}_{(x_{i},y_{i})\sim\mathcal{D}}\left[\mathcal{R}_{q}(x_{i},y_{i})-\beta\log{\beta}-1-\beta\log{\frac{\pi_{\theta}(y_{i}\mid x_{i})}{\pi_{ref}(y_{i}\mid x_{i})}}\right]^{2}\vskip-4.79993pt$